# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

### v4

@@ Abstract
With the advancement of Large Language Models (LLMs), more and more researchers apply LLMs for Named Entity Recognition (NER) methods, bringing vitality to this classical Natural Language Processing task. However, existing datasets are designed for traditional machine learning methods, inadequate for LLM-based methods in terms of corpus selection, entity categorization, and design logic. This limitation leads to less effective evaluation and model fine-tuning. To address this issue, we propose DynamicNER, the first NER dataset specifically designed for LLMs and with dynamic categorization, transcending the limitations of fixed categorization in existing datasets. It is also multi-lingual and multi-granular, covering 8 languages and 155 entity types, with corpus spanning multiple specialized domains. Furthermore, in response to the limitations demonstrated by existing LLM-based methods during DynamicNER testing, we develop CascadeNER, a novel NER method based on a two-stage strategy and lightweight LLMs, addressing the problems in current methods. Experiments show that DynamicNER is an effective benchmark for LLM-based NER methods, and CascadeNER outperforms existing methods with fewer computational resources. Our work is openly accessible.
@@
Hanjun Luothanks: Zhejiang University, hanjun.21@intl.zju.edu.cn 11footnotemark: 1, Yingbin Jinthanks: Zhejiang University, 3210104412@zju.edu.cn 22footnotemark: 2, Xinfeng Lithanks: Nanyang Technological University, xinfengli@zju.edu.cn 33footnotemark: 3, Xuecheng Liuthanks: Zhejiang University, xuecheng.21@intl.zju.edu.cn 44footnotemark: 4, Ruizhe Chenthanks: Zhejiang University, ruizhec.21@intl.zju.edu.cn 55footnotemark: 5, Tong Shangthanks: Southeast University, 213211756@seu.edu.cn 66footnotemark: 6, Kun Wangthanks: Nanyang Technological University, wk520529@mail.ustc.edu.cn 77footnotemark: 7, Qingsong Wenthanks: Squirrel AI, qingsongedu@gmail.com 88footnotemark: 8, Zuozhu Liuthanks: Zhejiang University, zuozhuliu@intl.zju.edu.cn, Corresponding author 99footnotemark: 9
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) brings revolutions to natural language processing (NLP) (Naveed et al., 2023). As a classical NLP task, Named Entity Recognition (NER) also witnesses significant method innovations driven by LLMs Xie et al. (2023). Existing LLM-based methods Shao et al. (2023); Li and Zhang (2023), leveraging LLMs’ generalization and contextual understanding capabilities, show superior performance compared to traditional machine learning (ML) methods in low-resource, multilingual, or few/zero-shot learning scenarios. This advantage is particularly crucial in AI for Science and Healthcare, where domain resources, especially non-English materials, are often limited. Thus, LLM-based methods gain attention from researchers in these fields Xiao et al. (2024).
Nevertheless, there are no existing NER datasets optimized for the characteristics of LLMs, limiting the effective training and evaluation of LLM-based methods. This limitation manifests primarily in three aspects. First, existing datasets employ static categorization with a fixed set of entity types, preventing effective evaluation of LLMs’ generalization capabilities in handling new entity types and varying granularities under few-shot/zero-shot scenarios. Secondly, the corpus in existing NER datasets focuses on short texts and isolated sentences, making them ineffective to evaluate the capabilities of LLMs in long-range context. Finally, while existing datasets separately address specialized corpus with corresponding entity types Kim et al. (2003); Liu et al. (2021), multi-grained classifications Ding et al. (2021), and multi-languages Malmasi et al. (2022), no dataset simultaneously incorporates three aspects, limiting comprehensive evaluation of LLM-based methods, which show potential across all these aspects. As a result, existing datasets lack the capability to effectively discriminate performance differences among LLM-based methods, fail to reflect their true potential and limitations, and impedes future improvements.
To address this issue, we develop DynamicNER, the first NER dataset optimized for LLM-based methods and the first NER dataset with dynamic categorization. It employs multiple strategies to dynamically adjust entity labels, type lists, and granularity levels during annotation, enhancing the evaluation of generalization ability of NER methods. We design cohesion and distribution balance metrics to enable evaluation and optimization during the categorization, with the entire procedure being algorithmically automated, ensuring reliability and reproducibility. This method addresses the limitations of existing datasets in training and evaluating few-shot/zero-shot learning ability. Besides, DynamicNER is a multilingual and multi-granular dataset, featuring 8 languages, 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, with entity types and corpora covering extensive professional content across domains including sciences, medicine, and art, providing an unprecedented level of coverage.
@@
Furthermore, our test on DynamicNER reveals significant limitations in existing methods, particularly when migrating to lightweight LLMs (LLMs with parameters from 1.5B to 7B) for local deployment. While methods with commercial models like GPT Brown (2020) achieve reported performance levels, this dependency introduces challenges regarding API costs and privacy risks. Using models via API are expensive for practical NER application. Privacy concern is also a significant issue Zhang et al. (2024); Das et al. (2024); Deng et al. (2025). Given the current absence of laws regarding LLM data uploading in many countries, users face difficulties in protecting their rights. Therefore, preventive measures to avoid unauthorized disclosure of information becomes essential.
@@
For this issue, we propose CascadeNER, a universal and multilingual NER framework that achieves comparable performance with lightweight LLMs to existing LLM-based methods requiring expensive commercial models. CascadeNER employs a two-stage strategy, redefines NER as two in-context text generation sub-tasks, extraction and classification. By model cascading Varshney and Baral (2022), CascadeNER assigns each sub-tasks to corresponding fine-tuned lightweight LLMs, decreasing task complexity and helping capture in-context dependencies. Leveraging the prior knowledge, CascadeNER can complete NER tasks in multiple languages in low-resource scenarios effectively.
@@
We evaluate a BERT-based Devlin et al. (2018) supervised method, two LLM-based methods, and our proposed CascadeNER on DynamicNER. The three LLM-based methods are tested on 3 difference models: Qwen2.5-1.5B, Qwen2.5-7B Bai et al. (2023), and GPT-4o Achiam et al. (2023). We also conduct evaluations of CascadeNER against existing methods on existing datasets. Results demonstrate that DynamicNER effectively evaluates the performance of LLM-based methods in low-resource and complex NER tasks, while CascadeNER outperforms existing LLM-based methods significantly with smaller models. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with a emphasis on multilingual and fine-grained scenarios.
Our core contributions are summarized as follows:
@@
  • We build DynamicNER, the first NER dataset optimized for LLM-based NER with a novel dynamic categorization system, which supports 8 languages, 155 entity types, and three-level granularity
@@
  • We propose CascadeNER, a universal and multilingual NER framework, which outperforms existing LLM-based methods with lightweight LLMs
@@
  • We conduct the first comprehensive comparison of LLM-based NER methods, discussing future directions of LLM-based NER
@@ 2 Related Works
Named Entity Recognition
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of train data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
NER Dataset
There have been a considerable number of NER datasets in various domains Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017). However, these existing datasets exhibit several limitations, making them unsuitable for LLM-based NER. Most previous multilingual NER datasets adopt coarse-grained classification, no longer meeting the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods. Table 1 presents a simple comparison between DynamicNER and existing multilingual or fine-grained datasets.
@@
DynamicNER 8 8 155 Multi Domain
@@ 3 DynamicNER Dataset
DynamicNER spans 8 languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of categorization, it is the first NER dataset with three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as shown in Figure 1. Like other NER datasets, DynamicNER is divided into train, dev, and test sets, and data volumes for different languages and parts shown in Appendix C. To develop DynamicNER, we first collect unlabeled corpus from Wikipedia. Then we manually extract sentences from corpora and annotate entities. After human annotation, we guide the dynamic categorization with category cohesion and distribution uniformity to automatically process base DynamicNER, and results in one base version and one dynamic version. Details are given in the following parts.
@@
Figure 1: The coarse-grained and medium-grained categories of DynamicNER. Detailed categories are provided in Appendix J.
Corpora Collection and Annotation
@@
Wikipedia provides multilingual, domain-specific corpora with clear hierarchical and indexing systems, serving as a rich resource for our research. We utilize legal Wikipedia-API to filter and download corpora across different languages and categories, followed by manual selection and annotation of sentences. We particularly focus on corpora containing long texts and complex contexts. After completing 50% of the annotation process of each language, we annotate corpora from categories related to underrepresented entity types to achieve a balanced entity type distribution. For instance, when the entities of "Algorithm" are significantly less than others, we use more corpora from Computer Science category. As a result, DynamicNER ensures balanced entity distribution, and includes rare entities and emerging fields which are not adequate in existing datasets, ensuring comprehensive coverage across diverse domains. A statement about annotators and ethical considerations is provided in Appendix I.
Dynamic Categorization
@@
The dynamic version improves model generalization and reduces overfitting risk by dynamically adjusting entity labels and corresponding entity type lists during annotation, including mixing types of different granularities, replacing types with synonyms, using type lists without irrelevant types, and merging certain types into miscellaneous/others, as shown in Figure 2. This method addresses the mismatch between existing datasets and few-shot/zero-shot training needs, better simulating real-world scenarios, and is particularly critical for evaluating methods relying on complex prompt designs (e.g., CoT). Unlike traditional few-shot learning, some LLM-based methods only use few-shot demos to help the model understand the task or format, without requiring knowledge of entity types. They can perform NER across different datasets with fixed few-shot demos, resembling zero-shot NER. Research shows this method is more effective than typical zero-shot NER Zhang et al. (2022). In methods that uses complex prompt designs like CoT to guide the reasoning, even few-shot CoT only conveys the CoT process rather than task-relevant knowledge, the performance of prompt-guide zero-shot CoT is significantly worse than few-shot CoT, making zero-shot restrictions inadequate for reflecting their true capability. However, in NER, models inevitably learn about entity types through few-shot demos, which limits generalization evaluations on fixed-category datasets. Our method significantly mitigates this limitation by varying entity types and lists, isolating the impact of prior type knowledge.
Figure 2: Examples of dynamic categorization. The red box contains the fine-grained categories in base version. Green boxes contains the lists modified with 4 methods. Blue boxes contains the entity type to each list.
Notably, as dynamic categorization is a subtractive method applied to a comprehensive classification system, this method highly relies on DynamicNER’s comprehensive categorization system, which includes 155 entity types. This method may not be suitable for all datasets.
Categorization Metrics
@@
Random dynamic categorization not only exhibits poor reproducibility and explainability, but may also lead to data quality degradation. For training, inappropriate categorization may result in inconsistent learning objectives and overfitting risks Ren et al. (2016). For evaluation, certain categories may experience imbalanced sample distribution and boundary ambiguity, reducing the comprehensiveness and consistency of evaluation Obeidat et al. (2019). Thus, we design four metrics to regulate the dynamic categorization: cohesion, normalized entropy, Gini coefficient, and variation coefficient. The definition and calculation methods are provided in Appendix A.
Figure 3: Pipeline of dynamic categorization. Blue boxes contains methods of each re-categorization. Yellow boxes contains used metrics of each evaluation.
Categorization Process
@@
The dynamic categorization process consists of 4 rounds of re-categorization, each sequentially corresponding to an adjustment method, and different metrics are employed in each round to guide the optimization. This hierarchical design enables each stage to focus on distinct data characteristics and optimization objectives, preventing interference between metrics while ensuring proper optimization direction, thus achieving a progressive optimization. We do not use all metrics in each evaluation, considering that certain metrics may have overlapping or conflicting effects at specific stages. For instance, normalized entropy and Gini coefficient both measure distribution uniformity, while improving cohesion may lead to more concentrated distribution and consequently lower entropy values. Figure 3 illustrates the metrics and methods corresponding to each round. Appendix B explains the reasons of metric selection for each round.

### v5

@@ Abstract
The advancements of Large Language Models (LLMs) have spurred a growing interest in their application to Named Entity Recognition (NER) methods. However, existing datasets are primarily designed for traditional machine learning methods and are inadequate for LLM-based methods, in terms of corpus selection and overall dataset design logic. Moreover, the prevalent fixed and relatively coarse-grained entity categorization in existing datasets fails to adequately assess the superior generalization and contextual understanding capabilities of LLM-based methods, thereby hindering a comprehensive demonstration of their broad application prospects. To address these limitations, we propose DynamicNER, the first NER dataset designed for LLM-based methods with dynamic categorization, introducing various entity types and entity type lists for the same entity in different context, leveraging the generalization of LLM-based NER better. The dataset is also multilingual and multi-granular, covering 8 languages and 155 entity types, with corpora spanning a diverse range of domains. Furthermore, we introduce CascadeNER, a novel NER method based on a two-stage strategy and lightweight LLMs, achieving higher accuracy on fine-grained tasks while requiring fewer computational resources. Experiments show that DynamicNER serves as a robust and effective benchmark for LLM-based NER methods. Furthermore, we also conduct analysis for traditional methods and LLM-based methods on our dataset. Our code and dataset are openly available at https://github.com/Astarojth/DynamicNER.
@@
[…]
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have transformed the landscape of NLP (Naveed et al., 2023). Among the impacted tasks, Named Entity Recognition (NER) has seen notable methodological shifts Xie et al. (2023). Leveraging LLMs’ strong generalization and contextual understanding capabilities, existing LLM-based methods Shao et al. (2023); Li and Zhang (2023) show superior performance compared to traditional machine learning (ML) methods Wang et al. (2020); Yan et al. (2021); Curran and Clark (2003) in low-resource, multilingual, or few-/zero-shot settings. As a result, LLM-based NER has garnered growing interest in these settings Xiao et al. (2024), offering a promising path toward more scalable and adaptable information retrieval.
Despite recent progress, there are no existing NER datasets specifically optimized for the characteristics of LLMs, thereby limiting both their effective evaluation and the development of optimized methods. Existing NER datasets employ static categorization with a fixed set of entity types, preventing the evaluation of LLMs’ ability to generalize to novel entity types and varying levels of granularity, especially in few-shot or zero-shot settings. Moreover, while some datasets address domain-specific corpora with specialized entity types Kim et al. (2003); Liu et al. (2021), others target multi-grained classifications Ding et al. (2021), or multilingualism Malmasi et al. (2022), no existing dataset simultaneously incorporates all three aspects. This fragmentation hinders comprehensive evaluation of LLM-based methods, which are particularly well-suited to handling such challenges. As a result, current datasets fall short in revealing performance differences between LLM-based methods, fail to capture their full potential and limitations, and ultimately impede the advancement of more effective NER solutions.
To address these gaps, we develop DynamicNER, the first NER dataset optimized for LLM-based methods and the first to support dynamic categorization. It employs multiple strategies to dynamically adjust entity labels, type lists, and granularity levels during annotation. This design simulates the complexity and uncertainty of entity types in real-world, general-purpose scenarios—a challenge that traditional ML methods with fixed entity categorization struggle to address, enabling a more rigorous evaluation of LLM-based NER methods’ ability to generalize across diverse and evolving scenarios. We introduce cohesion and distribution balance metrics to guide the evaluation and optimization of the annotation process. The entire procedure is algorithmically automated and openly available, ensuring both reliability and reproducibility.
In addition, DynamicNER is a multilingual and multi-granular dataset, featuring 8 languages, 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types. The multilingual and fine-grained nature of DynamicNER not only provides necessary support for the dynamic categorization, but also places higher demands on the NER methods. Many existing benchmarks are no longer difficult enough to reveal the performance boundaries of advanced methods, and the challenging DynamicNER will help future researchers to explore performance frontiers. Its entity types and corpora span a wide range of professional domains, including science, computer engineering, medicine, history, and arts. These features offer an unprecedented level of semantic and linguistic coverage for complex NER evaluation.
@@
Furthermore, our evaluation on DynamicNER reveals significant limitations in existing LLM-based methods, particularly when migrating to lightweight LLMs (models with 1.5B to 7B parameters) for local deployment. While approaches leveraging commercial models like ChatGPT Brown (2020) achieve high performance, this reliance introduces practical challenges related to API costs and privacy risks. API-based usage is often prohibitively expensive for real-world NER applications, and privacy remains a critical concern Zhang et al. (2024); Das et al. (2024); Deng et al. (2025).
@@
To address this issue, we propose CascadeNER, a universal and multilingual NER framework that achieves competitive performance with lightweight LLMs, comparable to existing LLM-based methods that rely on costly commercial models. CascadeNER employs a two-stage strategy by dividing NER as two in-context text generation sub-tasks, extraction and classification, instead of treating it as a traditional sequential labeling task. To reduce task complexity and better capture in-context dependencies, CascadeNER assigns each stage of the NER process–extraction and classification–to separate fine-tuned lightweight LLMs within a model cascading framework Varshney and Baral (2022). This modular architecture, combined with the integration of prior knowledge, enables effective multilingual performance in low-resource settings.
@@
We evaluate a BERT-based Devlin et al. (2018) supervised method, two LLM-based methods, and our proposed CascadeNER on DynamicNER. We also conduct evaluations of CascadeNER against existing methods on existing datasets. Results demonstrate that DynamicNER effectively evaluates the performance of LLM-based methods in low-resource and complex NER tasks, while CascadeNER outperforms existing LLM-based methods significantly with smaller models. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with an emphasis on multilingual and fine-grained scenarios.
Our contributions are summarized as follows:
@@
  • We develop DynamicNER, the first NER dataset optimized for LLM-based NER methods, featuring a novel dynamic categorization system. The dataset which supports 8 languages, 155 entity types, and 3 levels of granularity, enabling comprehensive evaluation across diverse linguistic and semantic settings.
@@
  • We propose CascadeNER, a universal NER framework, which outperforms existing LLM-based methods using only lightweight LLMs and a two-stage strategy.
@@
  • We conduct a comprehensive evaluation of LLM-based NER methods on our challenging dataset and identify key challenges and future directions for the field.
@@ 2 Related Works
Named Entity Recognition.
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of training data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
@@
DynamicNER (ours) 8 8 155 Multi Domain
@@
NER Datasets.
There have been a considerable number of NER datasets in various domains Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017); Katz et al. (2023). However, these existing datasets exhibit several limitations, making them unsuitable for LLM-based NER. Most previous multilingual NER datasets adopt coarse-grained classification, no longer meeting the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods. Table 1 presents a simple comparison between DynamicNER and existing multilingual or fine-grained datasets.
@@ 3 DynamicNER Dataset
Figure 1: The coarse-grained and medium-grained categories of DynamicNER. Detailed categories are provided in Appendix P.
DynamicNER spans 8 languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of categorization, it is the first NER dataset with three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as shown in Figure 1. Like other NER datasets, DynamicNER is divided into train, dev, and test sets. Data volumes for different languages and parts shown in Appendix D. To develop DynamicNER, we first collect unlabeled corpus from Wikipedia. Then we manually extract sentences from corpora and conduct comprehensive manual annotation on the collected sentences, strictly following our 3-level, 155-type fine-grained entity schema. This initial annotated result, containing all entity types without any automated adjustments, is what we define as the Base Version of DynamicNER. It represents our most complete, static knowledge base. Subsequently, using the Base Version as input, we apply the automated dynamic categorization according to four metrics, generating the Dynamic Version. To ensure reproducibility and support future research on dynamic categorization, we provide the full code for the pipeline, including all settings and parameters in our repository.
Corpora Collection and Annotation.
@@
Wikipedia provides multilingual, domain-specific corpora with clear hierarchical and indexing systems, serving as a rich resource for our research. We utilize legal Wikipedia-API to filter and download corpora across different languages and categories, followed by manual selection and annotation of sentences. We particularly focus on corpora containing long texts and complex contexts, with a special emphasis on encyclopedic content from professional domains such as science, medical, arts, engineering, and law, to ensure DynamicNER’s coverage of specialized domains. In addition, we have incorporated multilingual corpora from the social media platforms X and Weibo to ensure that our dataset also includes a substantial portion of colloquial language. After completing 50% of the annotation process of each language, we annotate corpora from categories related to underrepresented entity types to achieve a balanced entity type distribution. For instance, when the entities of "Algorithm" are significantly less than others, we use more corpora from Computer Science category. Thus, DynamicNER ensures balanced entity distribution, and includes rare entities and emerging fields that are inadequate in existing datasets, ensuring comprehensive coverage across diverse domains. Details about the manual annotation process is provided in Appendix J.
Figure 2: Examples of dynamic categorization. Details of the four strategies are provided in Appendix A.
Dynamic Categorization.
@@
The dynamic version improves model generalization and reduces overfitting risk by dynamically adjusting entity labels and corresponding entity type lists during annotation, including ➊ mixing types of different granularities, ➋ replacing types with synonyms, ➌ using type lists without irrelevant types, and ➍ merging certain types into miscellaneous/others, as shown in Figure 2. This method addresses the mismatch between existing datasets and few-shot/zero-shot training needs, better simulating real-world scenarios, and is particularly critical for evaluating methods relying on complex prompt designs (e.g., CoT). Unlike traditional few-shot learning, some LLM-based methods only use few-shot demos to help the model understand the task or format, without requiring knowledge of entity types. They can perform NER across different datasets with fixed few-shot demos, resembling zero-shot NER. Research shows this method is more effective than typical zero-shot NER Zhang et al. (2022). In methods that uses complex prompt designs like CoT to guide the reasoning, even few-shot CoT only conveys the CoT process rather than task-relevant knowledge, the performance of prompt-guided zero-shot CoT is significantly worse than few-shot CoT, making zero-shot restrictions inadequate for reflecting their true capability. However, in NER, models inevitably learn about entity types through few-shot demos, which limits generalization evaluations on fixed-category datasets. Our method significantly mitigates this limitation by varying entity types and lists, isolating the impact of prior type knowledge. Notably, as dynamic categorization is a subtractive method applied to a comprehensive and detailed categorization system, this method relies heavily on DynamicNER’s fine-grained taxonomy, which includes 3 levels of granularity and 155 entity types. Consequently, this method may not be suitable for all datasets. Based on our empirical findings, we recommend that future researchers adopting this categorization method use datasets that feature at least two levels of granularity and over 20 fine-grained entity types.
Categorization Metrics.
@@
Random dynamic categorization not only exhibits poor reproducibility and explainability, but may also lead to data quality degradation. For training, inappropriate categorization may result in inconsistent learning objectives and overfitting risks Ren et al. (2016). For evaluation, certain categories may experience imbalanced sample distribution and boundary ambiguity, reducing the comprehensiveness and consistency of evaluation Obeidat et al. (2019). Thus, we design four metrics to regulate the dynamic categorization: cohesion, normalized entropy, Gini coefficient, and variation coefficient. The definition and calculation methods are provided in Appendix B.
Categorization Process.
@@
The dynamic categorization process consists of 4 rounds of re-categorization, each sequentially corresponding to an adjustment method, and different metrics are employed in each round to guide the optimization. This hierarchical design enables each stage to focus on distinct data characteristics and optimization objectives, preventing interference between metrics while ensuring proper optimization direction, thus achieving a progressive optimization. We do not use all metrics in each evaluation, considering that certain metrics may have overlapping or conflicting effects at specific stages. For instance, normalized entropy and Gini coefficient both measure distribution uniformity, while improving cohesion may lead to more concentrated distribution and consequently lower entropy values. Figure 4 illustrates the metrics and methods corresponding to each round. Appendix C explains the reasons of metric selection for each round.
Figure 4: Pipeline of dynamic categorization.